### v1

# Crowd-SAM: SAM as a Smart Annotator for Object Detection in Crowded Scenes

Abstract
In computer vision, object detection is an important task that finds its application in many scenarios. However, obtaining extensive labels can be challenging, especially in crowded scenes. Recently, the Segment Anything Model (SAM) has been proposed as a powerful zero-shot segmenter, offering a novel approach to instance segmentation tasks. However, the accuracy and efficiency of SAM and its variants are often compromised when handling objects in crowded and occluded scenes. In this paper, we introduce Crowd-SAM, a SAM-based framework designed to enhance SAM’s performance in crowded and occluded scenes with the cost of few learnable parameters and minimal labeled images. We introduce an efficient prompt sampler (EPS) and a part-whole discrimination network (PWD-Net), enhancing mask selection and accuracy in crowded scenes. Despite its simplicity, Crowd-SAM rivals state-of-the-art (SOTA) fully-supervised object detection methods on several benchmarks including CrowdHuman and CityPersons. Our code is available at https://github.com/FelixCaae/CrowdSAM.
Keywords:
Detection in crowded scenes Few-shot learning Segment anything model
1 Introduction
Object detection in crowded scenes has emerged as a challenging task in computer vision, widely applied in areas such as traffic management, autonomous driving, and surveillance. The primary focus lies in identifying and locating densely packed common objects like pedestrians, where occlusion and clustering present significant challenges. Significant progress has been made in recent years, including two-stage methods[43, 57] and query-based methods[21, 60, 8]. However, these methods follow a supervised manner and necessitate extensive labeled training samples, incurring a considerable annotation cost of approximately 42.4 seconds per object[36] The density and complexity of crowded scenes further aggravate the annotation burden  [^1] On average, an image contains approximately 22 objects in CrowdHuman [35] and 7 in MS COCO [23].
The high cost of collecting object annotations drives the exploration of alternative approaches such as few-shot learning[42, 37, 32], weakly supervised learning [39, 50], semi-supervised learning [27, 47, 40, 28], and unsupervised learning[24, 6, 46, 1, 20]. The best-performing methods, i.e. semi-supervised object detection (SSOD), leverage the unlabeled data for training and achieve great success on common benchmarks such as PASCAL VOC[14] and COCO[23]. However, SSOD introduces extra complexity such as complicated augmentations and online pseudo-labeling.
In recent times, prompt-based segmentation models have attracted much attention due to their flexibility and scalability. Particularly, the Segment-Anything Model (SAM) [19] can effectively and accurately predict the masks of regions specified by prompts, in the any form of points, boxes, masks, or text descriptions. Recognizing its exceptional potential, researchers have made significant efforts to adapt it for various vision tasks such as medical image recognition[29], remote-sensing[4], industrial defect detection[49], etc.
Despite the huge progress following SAM  [17, 44, 51], most of the current work neglects the potential of applying SAM for object detection tasks in crowded scenes. In this paper, we investigate the potential of applying SAM for detection in crowded scenes with such motivations First, SAM is pre-trained on a very large dataset i.e. SA-1B that contains most of the common objects. It would be reasonable to utilize the learned knowledge to ease the difficulty of labeling massive data and training a brand-new detector. Second, SAM demonstrates robust segmentation ability in handling complicated scenes characterized by occlusion and clustered objects that could be difficult for an object detector trained from scratch.
In this paper, we propose Crowd-SAM, a smart annotator powered by SAM for object detection in crowded scenes. As depicted in  Fig. 1, we introduce a self-promoting approach based on DINOv2 to alleviate the cost of human prompting. Our method employs dense grids equipped with an Efficient Prompt Sampler (EPS) to cover as many objects as possible at a moderate cost. To distinguish the masks from multiple outputs precisely in occluded scenes, we propose a mask selection module, termed Part-Whole Discrimination Network (PWD-Net) that learns to distinguish the output with the highest quality. With a lightweight design and fast training schedule, it achieves considerable performance on benchmarks including CrowdHuman[35] and CityPersons[55].
Our contributions can be summarized as follows:
  1. 1.
    We introduce Crowd-SAM - a self-prompted segmentation method for labeling crowded datasets that can produce precise and accurate results with only a few examples.
  2. 2.
    We design two novel components of Crowd-SAM, i.e. EPS and PWD-Net which can effectively unleash SAM’s ability on crowded scenes.
  3. 3.
    We conduct comprehensive experiments on two benchmarks to demonstrate the effectiveness and generalizable nature of Crowd-SAM.
Figure 1: Pipeline comparison between SAM and Crowd-SAM. Crowd-SAM only requires a few labeled images and can automatically recognize target objects.
2 Related Work
Object detection. General object detection aims to identify and locate objects and is mainly divided into two categories: one-stage detectors and two-stage detectors. One-stage detectors predict bounding boxes and class scores directly from image features[33, 25, 22], while two-stage detectors first generate region proposals and then classify and refine these proposals [34, 10, 9]. Recently, end-to-end object detectors such as DETR[2, 61, 52] deprecate the hand-crafted modules such as non-maximum suppression (NMS) by adopting one-to-one matching in the training phase, showing great potential in a wide variety of areas.
However, applying these detectors directly to pedestrian detection tasks will lead to performance degradation due to the fact that pedestrians are often crowded and occluded. Early work[30] proposes integrating extra features into a pedestrian detector to explore the low-level visual cues, while follow-up methods [5, 53] attempt to utilize the head areas for better representation learning. In [53], an anchor is associated with two targets, the whole body and the head part, to achieve a more robust detector from joint training. Alternative methods focus on the design of loss functions to improve the training process. For example, RepLoss[43] encourages the prediction consistency of the same target while repels ones from different targets. Recently, Zheng et al. [60] model the relation of queries to improve DETR-based detectors in crowded scenes and achieves remarkable success. Although these works have pushed the boundaries of object detection in crowded scenes to a new stage, they all rely on a large number of labeled samples for training, which is labor-intensive. This limitation inspires us to develop label-efficient detectors and automatic annotation tools, with the help of SAM.
Few-shot object detection (FSOD). The task aims to detect objects of novel classes from limited samples. FSOD methods can be generally categorized into meta-learning based [48, 15] and fine-tuning based methods[42, 37, 32]. Meta-RCNN[48] processes the query and support images in parallel via a siamese network. The query’s region of interest (ROI) features are fused with the class prototypes to effectively transfer knowledge learned from the support set. TFA[42] proposes a simple two-stage fine-tuning method that only fine-tunes the detector’s last layers. FSCE[37] introduces a supervised contrastive loss in the fine-tune stage to mitigate misclassification issues. De-FRCN[32] stops the gradient from the RPN and scales the gradient from the R-CNN, followed by a Prototypical Calibration Block to refine the classification scores.
Segment Anything Models. SAM[19], proposed as a visual foundation model for the segmentation task, is trained on the SA-1B dataset using a semi-supervised learning paradigm. Its exposure to this vast repository of training samples renders it a highly capable class-agnostic model, effectively handling a wide range of objects in the world. Despite its impressive performance in solving segmentation tasks, it suffers from several issues like domain shift, inefficiency, class-agnostic design, etc. HQ-SAM [17] is proposed to improve its segmentation quality by learning a lightweight adapter. Fast-SAM [59] and Mobile-SAM [51] focus on fastening the inference speed of SAM by knowledge distillation. RSprompt[4] enables SAM to generate semantically distinct segmentation results for remote sensing images by generating appropriate prompts. Med-SA[45] proposes a Space-Depth Transpose method to adapt 2D SAM to 3D medical images and a Hyper-Prompting Adapter to achieve prompt-conditioned adaptation. Unfortunately, these approaches necessitate a considerable amount of labeled data for effective adaptation, rendering them impractical for crowded scenes where annotation costs are prohibitive. Different from them, Per-SAM[54] and Matcher [26] teaches SAM to recognize specified objects with only one or few instances by extracting training-free similarity priors. SAPNet [44] proposes a weakly-supervised pipeline for instance segmentation. Although these approaches reduce the data requirement, they still lag behind the demands of crowded scenes, such as pedestrian detection, particularly in occlusion scenes.
Figure 2: The pipeline of Crowd-SAM shows the interaction between different modules. DINO encoder and SAM are frozen in the training process. * represents the parameters that are shared. For simplicity, the projection adapter of DINO is dismissed.
3 Our Approach
3.1 Preliminaries
SAM [19] is a promising segmentation model that comprises three main components: (a) an image encoder responsible for feature extraction; (b) a prompt encoder designed for encoding the geometric prompts provided by the users; and (c) a lightweight mask decoder that predicts the masks conditioned on the given prompts. Leveraging extensive training data, SAM demonstrates impressive zero-shot segmentation performance across various benchmarks. In particular, SAM utilizes points and boxes as prompts to specify the interested regions.
DINO [3] represents a family of self-supervised vision transformers [7] designed for general-purpose applications. During its training, DINO employs a self-distillation approach akin to BYOL [11], fostering the learning of robust feature representations. DINOv2 [31] enhances the foundation of DINO by integrating additional pre-training tasks, improving its scalability and stability, especially in large models like ViT-H with 1 billion parameters. Thanks to its enhancement, DINOv2 demonstrates strong feature representation ability, especially for the task of semantic segmentation.
3.2 Problem Definition and Framework
As shown in  Fig. 1, our goal is to detect pedestrians in crowded scenes with few annotated data. We formulate this problem as a one-class few-shot detection task. A common few-shot approach is to split the data into the base split and the novel split. Differently, we directly use the data of the target class to train our model, since the foundation models have already been trained on massive data. In particular, we use segmentation masks as intermediate results which can be easily converted to bounding boxes. During the training and evaluation process, only box annotations are provided.
Naive study on SAM auto generator. Before we introduce our method, we take a deeper look into how the prompt number affects SAM’s performance in crowded scenes. To this end, we conduct several naive studies on CrowdHuman[35] with SAM’s auto generator, which utilizes grid points to search every region.  Tab. 1 conveys three key observations: (1) dense grids are necessary for crowded scenes; (2) the ambiguity of point prompts and class-agnostic prompts distribution incur many false positives (FPs); (3) decoding time is a non-negligible burden when grid size is large. It is concluded that the dense prompts and FP removal are key aspects in designing SAM-based methods for detection/segmentation tasks in crowded scenes.
N 16 32 64 128 Oracle
Recall 33.6 58.0 63.4 76.0 91.4
avg. FP 51 112 227 485 -
T(s) 0.059 0.22 0.83 3.2 0.045
Table 1: Comparisons of different grid sizes adopted by SAM generator on Crowd-Human[35]. Oracle model derives the prompts by computing the center of the ground truth boxes. T is the decoding time collected on a 3090 Ti GPU card.
Framework. Inspired by the above studies, we equip SAM [19] with several proper components to achieve an accurate and efficient annotation framework, as illustrated in  Fig. 2. To accurately locate pedestrians, we employ the foundational model DINOv2[31] to predict a semantic heat map, a task that can be formulated with a simple binary classifier. To discriminate the output masks that is a mixture of correct masks, backgrounds, and part-level masks, we adopt Part-Whole Discrimination Network that takes as input both the learned tokens from SAM and the semantic-rich tokens from DINOv2 to re-evaluate all the outputs. Finally, to handle the redundancy brought by the use of dense grids, we propose an efficient prompt sampler (EPS) to decode the masks at a moderate cost.
We introduce the details of our methods in the following sections.
3.3 Class-specific Prompts Generation
Generating a unique point prompt for each pedestrian in the dense scenes is de facto a non-trivial problem. Thus, we take a step back and study how to detect pedestrians with multiple prompts associated with one person, and then use the proper post-processing methods to remove duplicates. To this end, we adopt a heat-map-based prompt generation pipeline that classifies the regions first and generates prompts from the positive regions.
For the input image $I\in\mathbb{R}^{H\times W\times 3}$, we use a pre-trained image encoder $E_{D}$ to extract semantic rich features first. To better transfer the pre-trained features to pedestrian segmentation, we add an MLP block after the final output layer and thus get the adapted features $F_{DINO}\in\mathbb{R}^{{H\over s}\times{W\over s}\times C}$, where $s$ is the patch size of DINOv2 [31] and $C$ is the output channel. Then, we use a segmentation head $\text{Head}_{CLS}$ to classify $F_{DINO}$ pixel by pixel, resulting in a heatmap ^$\hat{\mathbb{H}}\in[0,1]^{{H\over s}\times{W\over s}}$ that indicates the locations of pedestrians as ^$\hat{\mathbb{H}}=\text{Head}_{CLS}(F_{DINO})$.
Given only the bounding box annotations $B\in[0,1]^{N\times 4}$, where $N$ is the number of targets, this binary segmentation head can be optimized with box-level supervision. However, the coarse boundaries would incur considerable points scattering in the background regions. To alleviate this issue, we use SAM [19] to obtain high-quality mask-level pseudo labels, which is illustrated in  Fig. 4, with ground truth (GT). The decoded masks are then merged into a single foreground mask $\mathbb{H}\in\{0,1\}^{256\times 256}$. We use dice loss for training the adapter and segmentation head with the generated pseudo masks as follows:
$^{\cal L}_{fg}=dice(f(\hat{\mathbb{H}}),\mathbb{H}),$ (1)
where $f$ is a up-sampling function that resizes $\mathbb{H}$ to $256\times 256$. During inference, we use a threshold $t$ for mask binarization, which is simply set 0.5 in our experiments. The binarized masks are mapped to point prompts $P_{G}$ which only contain those in positive regions.
3.4 Semantic-guided Mask Prediction
Given the proposals generated in  Sec. 3.3, this section aims to decode the dense prompts efficiently and discriminate the generated masks accurately. As depicted in  Fig. 3, each instance contains many prompts due to the density of grids. Supposing only one in-position prompt is required for mask prediction for SAM, decoding all the prompts would lead to not only a waste of computation but also cause more FPs due to some badly located prompts.
Figure 3: Illustration of EPS. The PWD-Net produce valid masks with a confidence threshold $T$. In each iteration, we prune prompts that fall inside valid masks (with a cross above).
Efficient Prompt Sampler. To address this challenge, we introduce the Efficient Prompt Sampler (EPS), which dynamically prunes prompts based on the confidence of decoded masks. This method is elaborated in  Algorithm 1. Beginning with the generated point prompts list $P_{G}$, EPS, in each iteration, samples a batch of prompts $P_{B}$ from $P_{G}$ using uniform random sampling. The sampled prompts $P_{B}$ are then appended to the output points list $P_{S}$. Subsequently, we employ SAM Generator with the batched prompts $P_{B}$ to produce masks $M$. Furthermore, we aggregate discriminant confidence scores $S$ from PWD-Net, which are employed to select valid masks $M^{\prime}$ using a score threshold $T$, by using $M^{\prime}=M[S>T]$. A detailed exposition of PWD-Net will be provided in the subsequent section. The valid masks represent regions that are believed to be well segmented, thus we remove the points that are already covered by any mask $m$ in $M^{\prime}$ from the prompts list $P_{G}$. The iteration halts when $P_{G}$ is empty. Furthermore, EPS establishes a stopping criterion with a parameter $K$, terminating the sampling process once the total sample count reaches this limit. This parameter is instrumental in managing the overall decoding cost.
Input: Feature extracted from image $I$ by image encoder of SAM: $F_{SAM}$; Generated prompt list: $P_{G}=\{p_{1},p_{2},...,p_{N}\}$; SAM Generator: $G$; Mask confidence threshold: $T$; Target sampled prompt list size: $K$;
Output: Sampled prompt list: $P_{S}$; Valid masks list: $M_{S}$.
1 $P_{S}\leftarrow\emptyset$;
2while $|P_{G}|>0$ and $|P_{S}|<K$ do
3       Sample a batch of prompts $P_{B}\subseteq P_{G}$ with uniform random sampling;
4      $P_{S}\leftarrow P_{S}\cup P_{B}$;
5      $P_{G}\leftarrow P_{G}\setminus P_{B}$;
6      Generate masks $M$ corresponding to $P_{B}$ by $G(F_{SAM},P_{B})$;
7      Select valid masks $M^{\prime}$ according to $T$ from $M$;
8      $M_{S}\leftarrow M_{S}\cup M^{\prime}$;
9      for $p\in P_{G}$ do
10             if $\exists m\in M^{\prime}\,\text{such that}\,p\in m$ then
11                   $P_{G}\leftarrow P_{G}\setminus\{p\}$
12             end if
13
14       end for
15
16 end while
Algorithm 1 Process of Efficient Prompt Sampler (EPS).
Part-Whole Discrimination Network. Given the raw masks predicted by SAM conditioned on the sampled batch of prompts $P_{B}$, we are inspired to design an automatic selector for choosing the best fitting mask. It should refine the outputs in two aspects: (i). help to refine the output IoU score according to the quality of related masks, if they are positive samples; (ii). help suppress the scores of samples that fall in the background regions. usion of those associated with the background.
Illustrated in  Fig. 2, for masks generated corresponding to $N$ prompts, we leverage the Mask Tokens and IoU Tokens within the mask decoder of SAM along with the sophisticated features extracted by the self-supervised pre-trained model DINOv2 [31]. ${\cal M}$ and ${\cal U}$ are responsible for mask decoding and IoU prediction in SAM Mask Decoder, respectively. Thus, we suppose that they contain shape-aware information, which is helpful in discriminating the mask.These components enable us to compute a discriminant confidence score $S$ for each specific prompt in a few-shot approach. Initially, the refined IoU score $S_{iou}$ is computed as follows:
$S_{iou}=\text{Head}_{par}(\text{Concat}(\text{Repeat}({\cal U}),{\cal M}))+\text{Head}_{IoU}({\cal U}),$ (2)
where $\text{Head}_{par}$ is a parallel IoU head, consisting of a MLP block, ${\cal U}\in\mathbb{R}^{N\times 1\times C}$ and ${\cal M}\in\mathbb{R}^{N\times 4\times C}$ denote the IoU Tokens and Mask Tokens respectively. Notice in  Eq. 2, the $S_{iou}$ is a sum of outputs from two individual heads, the parallel adapter head $\text{Head}_{par}$ and the original $\text{Head}_{IoU}$. We freeze the parameters of $\text{Head}_{IoU}$ to avoid overfitting in few-shot learning. Moreover, since these two tokens, ${\cal M}$ and ${\cal U}$ have different shapes, we repeat ${\cal U}$ by 4 times and then concat these two tokens in the channel dimension. The refined score $S_{iou}\in\mathbb{R}^{N\times 4}$ encapsulates the quality of the generated masks by assessing the texture-aware feature from ${\cal M}$ and ${\cal U}$.
Secondly, by harnessing the semantic feature embedded within the self-supervised pre-trained model DINOv2, along with the mask data $\hat{M}$, we calculate the discriminant score $S_{cls}$ :
$S_{cls}=\sigma({\text{Head}_{CLS}({\cal O})}),{\cal O}=\text{Pool}(d(\text{Softmax}(\hat{M}))\circ F_{DINO}).$ (3)
Here, we use ${\cal O}\in\mathbb{R}^{N\times C}$ to denote the extracted Semantic Token, $\hat{M}\in\mathbb{R}^{N\times H\times W\times 4}$ and $F_{DINO}\in\mathbb{R}^{N\times h\times w\times C}$ to represent the masks generated by SAM[29] and the features extracted by the DINOv2[31] model, respectively. We denote $d$ as a down-scale function that resizes the mask to $h\times w$ to be consistent with $F_{DINO}$. Pool is a global pooling function that conducts mean pooling on the x-axis and y-axis.The discriminant score $S_{cls}\in\mathbb{R}^{N\times 4}$ predicts whether masks belong to the foreground or background. The $\text{Head}_{CLS}$ shares the same parameters with the binary classifier introduced in  Sec. 3.3. Finally, we calculate a joint score of discrimination and estimate quality for masks by simply multiplying the two scores together: $S=S_{iou}\cdot S_{cls}$.
During training, prompts sampled within real masks are utilized as input. Regarding prompts within the foreground, the discriminant confidence score should aim to accurately predict the Intersection over Union (IoU) of generated and real masks. Conversely, for prompts within the background, the score should ideally output 00, indicating their association with the background. Hence, the loss function for this aspect is formulated as follows:
$s^{i}_{target}=\begin{cases}mIoU(m^{i},m^{i}_{GT}),&m^{i}_{GT}\in M^{bg}_{GT},\\ 0,&m^{i}_{GT}\in M^{fg}_{GT},\end{cases}$ (4)
${\cal L}_{iou}=\text{mse}(S,S_{target})$ (5)
Here, $s^{i}_{target}$ denotes the target score of the mask $m^{i}$ generated by the $i^{th}$ prompt, and $S_{target}=\{s^{i}_{target}\}_{i=1}^{N};M=\{m^{i}\}_{i=1}^{N};M_{GT}=\{m^{i}_{GT}\}_{i=1}^{N}=M_{GT}^{bg}\cup M_{GT}^{fg}$.
3.5 Training and Inference
The total training loss of our framework is the combination of  Eq. 1 and Eq. 5:
${\cal L}={\cal L}_{fg}+{\cal L}_{iou}$ (6)
During inference, we select the mask with the highest confidence score $S$ among the $4$ masks as the output of PWD-Net, which we denote as $M\in\mathbb{R}^{N\times H\times W}$ and $S_{o}\in\mathbb{R}^{N}$. We adopt a window cropping strategy for better performance on small objects as [19]. This strategy slices the whole image into overlapping crops where each crop is processed individually. The final results are merged from the outputs of each crop. We apply non-maximum suppression on the final outputs to remove duplicate proposals.
4 Experiments
Method Backbone #Shot AP MR-2 Recall Secs/Img
Fully supervised object detectors
ATSS [56] ResNet-50 [13] Full 80.3 59.7 86.1 0.051
FCOS [41] ResNet-50 [13] Full 76.3 65.5 82.6 0.045
Iter-SRCNN [60] ResNet-50 [13] Full 85.9 58.3 93.3 0.25
DINO [52] ResNet-50[13] Full 86.7 57.6 94.5 0.27
Few-shot object detectors
TFA [42] ResNet-101 [13] 10 46.9 84.3 57.9 0.067
FSCE [37] ResNet-101 [13] 10 43.0 84.7 50.0 0.072
De-FRCN [32] ResNet-101 [13] 10 46.4 85.9 65.5 0.072
SAM-based approaches
SAM [19] ViT-L [7] 0 - - 65.6 1.3
SAM* [19] ViT-L [7] 0 - - 79.6 6.7
Matcher [32] ViT-L [7] 1 8.0 88.9 23.9 22.0
Crowd-SAM ViT-L[7] 10 71.4 75.1 83.9 1.7
Crowd-SAM* ViT-L[7] 10 78.4 74.8 85.6 8.1
Table 2: Results on CrowdHuman[35] val set. All SAM-based methods adopt ViT-L[7] as the pre-trained backbone. SRCNN represents Sparse R-CNN[38] which is a baseline to [60]. * represents using the multi-crop trick.
Datasets. Following [60], we adopt CrowdHuman[35] as the benchmark to conduct main experiments and ablation studies. CrowdHuman[35] collects and annotates images containing crowded persons from the internet. It contains 15,000, 4,370, and 5,000 images for training, validation, and testing, respectively. We also test our method on CityPersons[55] for a realistic urban-scene evaluation. Additionally, to further investigate our method, we utilize OCC-Human [58] for evaluation, which is specially designed for occluded persons. For these pedestrian datasets, we utilize visible annotations (only including visible areas of an object) for training and evaluation. To validate the extensibility of Crowd-SAM, we also devise a multi-class version of Crowd-SAM by adding a multi-class classifier. We utilize 0.1 % percent of the COCO [23] trainval set for training and COCO val set for validation. We also validate our methods on an occluded subset of COCO, COCO-OCC [16] that is extracted by selecting images whose objects have a high overlapping ratio.
Implementation details. We utilize SAM (ViT-L) [19] and DINOv2 (ViT-L)[31] as the base models for all experiments. In the fine-tuning stage, their parameters are all frozen to avoid over-fitting. Instead of real GT, we use pseudo masks generated by SAM (ViT-L) as $S_{target}$ to supervise the learning of PWD-Net in Eq. 5. These generated pseudo labels are of high quality as shown in Fig. 4(c). We randomly pick points from the pseudo masks as positive training samples and points from the background part as negative training samples. In the training process, we use Adam[18] with a learning rate of $10^{-5}$, a weight decay of $10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.99$ for optimization. We train our module for 2000 iterations with a batch size of 1, which can be done on a single GTX 3090 Ti GPU in several minutes. For more details, please refer to the appendix.
Evaluation metrics. Following [60], we use AP with IoU threshold 0.5, MR-2, and Recall for our evaluation metrics. Generally, a higher AP, Recall, and lower MR-2 indicates a better performance.
4.1 Experimental Results on Pedestrian Detection
For a fair comparison, we re-implement the involved methods[56, 41, 60, 52, 32] with a 2×2$\times$2 × schedule using the visible annotations in the CrowdHuman [35] and CityPersons [55] dataset.
Main results. We compare Crowd-SAM with most related methods including fully supervised object detectors[41, 56, 52, 60], few shot object detectors[32] and SAM-based approaches[26, 19] for comprehensive understanding. Notice we use the visible annotation for which derives different results from those on full box annotation.
As shown in  Tab. 2, with only 10 labeled images, our method achieves comparable performance with full-supervised object detectors whose best result is 86.7% AP, achieved by DINO[52]. Particularly, Crowd-SAM outperforms an advanced anchor-free detector FCOS[41] by 2.1% AP. These results indicate that with proper adaptation techniques, SAM can achieve very competitive performance on complex pedestrian detection datasets like CrowdHuman. On the other hand, Crowd-SAM achieves SOTA performance on few-shot detection settings and outperforms all few-shot object detectors with a significant margin. DeFRCN [32] is a well-established few-shot object detector equipped with ResNet-101[13] that achieves 46.4% in AP. Notably, our method outperforms it by 32%, indicating the superiority of Crowd-SAM in the few-shot detection setting. The qualitative comparisons between ours and De-FRCN are depicted in  Fig. 4. For SAM-based approaches, our method largely leads the SAM baseline by 6% (with multi-crop) and 18.3% (w/o multi-crop) in the recall. Crowd-SAM is also superior to other methods such as Matcher[26].
Results on OccHuman and CityPersons. To investigate the performance of Crowd-SAM in occluded scenes, we compare it with an advanced supervised method Pose2Seg [58], and the results are shown in Tab. 3. It is noteworthy that Pose2Seg is a full supervised detector while Crowd-SAM is a few-shot detector. As can be seen from the results, Crowd-SAM leads Pose2Seg by 9.2% AP which demonstrates its robustness in occluded scenes. Also, we apply our method to a not-so-crowded but more realistic urban dataset, CityPersons [55], as reported in Tab. 4. Crowd-SAM outperforms TFA [42] by 2.7% AP, FSCE [37] by 1.1% AP, and De-FRCN [32] by 7.8% AP. In conclusion, our method remains competitive compared to advanced few-shot object detectors, even though it is not specifically designed for sparse scenes.
These results illustrate that our method has fully unleashed the power of vision foundation model models, i.e. SAM and DINO, in all of the crowded, occluded, and urban scenes.
Method Backbone AP APM APH
Mask R-CNN [12] R50-FPN 16.3 19.4 11.3
Pose2Seg [58] R50-FPN 22.2 26.1 15.0
Crowd-SAM ViT-L 31.4 26.5 17.7
Table 3: Comparative results on OccHuman [58] val. APM and APH represent AP in moderate and hard cases according to occlusion ratio, respectively.
Method # Shot AP MR-2
FCOS[41] Full 58.8 30.0
ATSS[56] Full 54.1 27.8
Iter-SRCNN[60] Full 57.9 31.0
TFA[42] 50 30.6 53.8
FSCE[37] 50 32.2 46.5
De-FRCN[32] 50 25.5 67.1
Crowd-SAM 50 33.3 31.7
Table 4: Comparative results on CityPersons[55] val.
Methods Backbone COCO-OCC COCO
AP AP50 AP AP50
Faster R-CNN [34] R-50-FPN 29.7 50.0 33.5 53.7
BCNet [16] R-50-FPN 31.7 51.1 34.6 54.4
Crowd-SAM ViT-L 20.6 31.5 22.0 33.7
Table 5: Comparative results on COCO [23] and COCO-OCC [16] val.
4.2 Experimental Results on Multi-class Object Detection
To further explore the extensibility of our methods in a more popular setting, we devise a multi-class Crowd-SAM. The multi-class version of Crowd-SAM is slightly different by replacing the binary classifier with a multi-class classifier. We validate multi-class Crowd-SAM on COCO [23], a widely adopted common object detection benchmark, and COCO-OCC [16] which is a split of COCO that mainly consists of images with high occlusion ratio.
We compare our method with two supervised detectors, Faster R-CNN [34] and BCNet [16], and report the results in Tab. 5. It is seen that our Crowd-SAM is comparable to supervised detectors in both datasets and drops only 1.4 APwhen comparing COCO and COCO-OCC. This minor drop indicates that our method is robust to the occlusion.
FG loc. EPS PWD-Net Multi-Crop AP MR-2 Recall
78.4 74.8 85.6
71.0 77.9 77.1
77.8 71.8 83.9
17.0 99.6 83.1
71.4 75.1 83.9
8.7 99.8 70.1
Table 6: Ablation studies on the main components in Crowd-SAM. FG location represents the use of binary classifier on DINOv2 features. The last line represents a SAM baseline with a standard $32\times 32$ grid as inputs.
4.3 Ablation Studies
Ablation on the modules. We conduct ablation studies on the key components of Crowd-SAM - foreground location, EPS, and PWD-Net - to validate their effectiveness. In Tab. 6, the performance on AP significantly drops by 7.4% and recall by 8.5% when the FG location is removed, indicating the importance of restricting foreground areas. As for EPS, once it is removed, the AP drops by 0.6%. We suppose EPS not only accelerates the sampling process of dense prompts but also helps focus on the difficult part of the images, which has ambiguous semantics. We also compare EPS with other batch iterators in  Tab. 7. Moreover, we find the PWD-Net is indispensable and when removed, the AP dramatically drops to only 17.0%. Finally, we point out that multi-cropping is a strong method to enhance the performance which contributes 7.4 % AP to the final performance. Overall, these results prove that all components are essential.
Full Random EPS
Grid K AP Recall AP Recall AP Recall
32 $\times$ 32 500 57.6 60.5 57.6 60.5 57.0 60.0
64 $\times$ 64 500 69.4 73.9 69.4 73.9 69.3 73.6
128 $\times$ 128 500 OOM 69.7 74.1 72.4 77.8
192 $\times$ 192 500 OOM 69.8 73.7 73.2 78.2
256 $\times$ 256 500 OOM 68.9 73.0 72.3 78.0
Table 7: Comparisons of different sampler. Full means using all prompts. OOM represents out-of-memory error which occurs when the GPU memory is all consumed.
Ablation on EPS. We conduct a comprehensive study on EPS by comparing it with several variants. We forward each image only once to avoid the side effects caused by multi-crop. We use a default sampler that iterates through all training samples and a random iterator with a halting threshold $K$ as counter-parts. As reported in  Tab. 7, AP and Recall increase with the grid size for all three samplers. However, the default sampler meets an out-of-memory error when the grid size reaches 128, preventing it from being adopted in this setting. As for the random sampler, its performance is constrained by K and a larger grid size than 64 only leads to limited improvement, e.g. 0.1% AP. On the contrary, our EPS can benefit from a much larger grid size and achieves a better performance especially when the grid size is large.
${\cal M}$ ${\cal U}$ ${\cal O}$ IoU head AP MR-2
$F$ 78.4 74.8
$T$ 75.6(-2.8) 80.4(+5.6)
$F$ 77.3(-1.1) 73.5(-1.3)
$F$ 76.3(-2.1) 74.8(-0.0)
$F$ 38.4(-40.0) 95.8(+21.0)
Table 8: Ablation studies on the design of PWD-Net. ${\cal M}$, ${\cal U}$ and ${\cal O}$ represent mask token, IoU token, and semantic token, respectively. In the column of the IoU head, $F$ means to freeze the original IoU head and train a parallel one and $T$ means tuning the original IoU head.
Ablation on PWD-Net. We compare PWD-Net with variants that replace some tokens with a full-zero placeholder. We also compare with two designs, tuning the IoU head directly or learning a parallel IoU head. As shown in  Tab. 8, all three tokens, i.e. Mask Token ${\cal M}$, IoU Token ${\cal U}$, and Semantic Token ${\cal O}$, contribute to the final result. Particularly, once ${\cal M}$ is removed, the AP drops by 40.0%, which is a catastrophic decline. This degradation suggests that the Mask Token contains the essential shape-aware feature that is essential for the part-whole discrimination task. Notably, the AP drops by 2.8% when we tune the pre-trained IoU Head, suggesting that it might overfit the few labeled images. We believe by freezing the SAM’s IoU head, the PWD-Net can benefit more from the shape-aware knowledge learned on massive segmentation data.
Figure 4: Qualitative comparisons between our proposed Crowd-SAM (a) and De-FRCN (b). We also plot the generated masks and GT boxes (c) and our prompts filtering results (d), where red points are kept and gray points are filtered.
5 Conclusion
This paper introduces Crowd-SAM, a SAM-based framework for object detection and segmentation in crowded scenes, designed to streamline the annotation process. For each image, Crowd-SAM generates dense prompts for high recall and uses EPS to prune redundant prompts. To achieve accurate detection in occlusion cases, Crowd-SAM employs PWD-Net that leverages several informative tokens to select the best-fitting mask. Combined with the proposed methods, Crowd-SAM achieves 78.4 AP on Crowd Human, comparable to full-supervised detectors, validating that object detection in crowded scenes can benefit from foundation models like SAM in data efficiency.
Acknowledgements
This work is partly supported by the National Natural Science Foundation of China (No. 62022011), the Research Program of State Key Laboratory of Software Development Environment, and the Fundamental Research Funds for the Central Universities.

### v2

@@
 Abstract
-In computer vision, object detection is an important task that finds its application in many scenarios.
-However, obtaining extensive labels can be challenging, especially in crowded scenes.
-Recently, the Segment Anything Model (SAM) has been proposed as a powerful zero-shot segmenter, offering a novel approach to instance segmentation tasks.
-However, the accuracy and efficiency of SAM and its variants are often compromised when handling objects in crowded and occluded scenes.
-In this paper, we introduce Crowd-SAM, a SAM-based framework designed to enhance SAM’s performance in crowded and occluded scenes with the cost of few learnable parameters and minimal labeled images.
-We introduce an efficient prompt sampler (EPS) and a part-whole discrimination network (PWD-Net), enhancing mask selection and accuracy in crowded scenes.
-Despite its simplicity, Crowd-SAM rivals state-of-the-art (SOTA) fully-supervised object detection methods on several benchmarks including CrowdHuman and CityPersons.
+Object detection is an important task that finds its application in a wide range of scenarios.
+Generally, it requires extensive labels for training, which is quite time-consuming, especially in crowded scenes.
+Recently, Segment Anything Model (SAM) has emerged as a powerful zero-shot segmenter, offering a novel approach to instance segmentation.
+However, the accuracy and efficiency of SAM and its variants are often compromised when handling objects in crowded scenes where occlusions often appear.
+In this paper, we propose Crowd-SAM, a SAM-based framework designed to enhance the performance of SAM in crowded scenes with the cost of few learnable parameters and minimal labeled images.
+We introduce an efficient prompt sampler (EPS) and a part-whole discrimination network (PWD-Net), facilitating mask selection and contributing to an improvement in accuracy in crowded scenes.
+Despite its simplicity, Crowd-SAM rivals state-of-the-art fully-supervised object detection methods on several benchmarks including CrowdHuman and CityPersons.
 Our code is available at https://github.com/FelixCaae/CrowdSAM.
@@
 1 Introduction
-Object detection in crowded scenes has emerged as a challenging task in computer vision, widely applied in areas such as traffic management, autonomous driving, and surveillance.
-The primary focus lies in identifying and locating densely packed common objects like pedestrians, where occlusion and clustering present significant challenges.
-Significant progress has been made in recent years, including two-stage methods[43, 57] and query-based methods[21, 60, 8].
-However, these methods follow a supervised manner and necessitate extensive labeled training samples, incurring a considerable annotation cost of approximately 42.4 seconds per object[36] The density and complexity of crowded scenes further aggravate the annotation burden  [^1] On average, an image contains approximately 22 objects in CrowdHuman [35] and 7 in MS COCO [23].
-The high cost of collecting object annotations drives the exploration of alternative approaches such as few-shot learning[42, 37, 32], weakly supervised learning [39, 50], semi-supervised learning [27, 47, 40, 28], and unsupervised learning[24, 6, 46, 1, 20].
-The best-performing methods, i.e.
-semi-supervised object detection (SSOD), leverage the unlabeled data for training and achieve great success on common benchmarks such as PASCAL VOC[14] and COCO[23].
-However, SSOD introduces extra complexity such as complicated augmentations and online pseudo-labeling.
-In recent times, prompt-based segmentation models have attracted much attention due to their flexibility and scalability.
-Particularly, the Segment-Anything Model (SAM) [19] can effectively and accurately predict the masks of regions specified by prompts, in the any form of points, boxes, masks, or text descriptions.
-Recognizing its exceptional potential, researchers have made significant efforts to adapt it for various vision tasks such as medical image recognition[29], remote-sensing[4], industrial defect detection[49], etc.
-Despite the huge progress following SAM  [17, 44, 51], most of the current work neglects the potential of applying SAM for object detection tasks in crowded scenes.
-In this paper, we investigate the potential of applying SAM for detection in crowded scenes with such motivations First, SAM is pre-trained on a very large dataset i.e.
-SA-1B that contains most of the common objects.
-It would be reasonable to utilize the learned knowledge to ease the difficulty of labeling massive data and training a brand-new detector.
-Second, SAM demonstrates robust segmentation ability in handling complicated scenes characterized by occlusion and clustered objects that could be difficult for an object detector trained from scratch.
-In this paper, we propose Crowd-SAM, a smart annotator powered by SAM for object detection in crowded scenes.
+Object detection in crowded scenes is a fundamental task in areas such as autonomous driving and video surveillance.
+The primary focus lies in identifying and locating densely packed common objects like pedestrians and vehicles, where occlusions present significant challenges.
+Great progress has been made in recent years, including two-stage methods [45, 60] and query-based methods [22, 62, 8].
+However, these methods mainly follow a supervised manner and necessitate extensive labeled training samples, incurring a considerable annotation cost of approximately 42.4 seconds per object [38] The density and complexity of crowded scenes further aggravate the annotation burden.
+[^1] On average, an image contains approximately 22 objects in CrowdHuman [37] and 7 in MS COCO [24]
+The high cost of collecting object annotations drives the exploration of alternatives such as few-shot learning [44, 39, 34], weakly supervised learning [41, 53], semi-supervised learning [29, 49, 42, 30], and unsupervised learning [26, 6, 48, 1, 21].
+The best-performing ones, i.e.
+semi-supervised object detection (SSOD), leverage both labeled and unlabeled data for training and achieve a big success on common benchmarks e.g.
+PASCAL VOC [15] and COCO [24].
+Unfortunately, SSOD introduces extra complexity such as complicated augmentations and online pseudo-labeling.
+Recently, prompt-based segmentation models have received increasing attention due to their flexibility and scalability.
+Particularly, Segment Anything Model (SAM) [20] show its high capability to effectively and accurately predict the masks of regions specified by prompts, in any form of points, boxes, masks, or text descriptions.
+Recognizing its exceptional potential, researchers have made many efforts to adapt it for various vision tasks such as medical image recognition [31], remote sensing analysis [4, 12], industrial defect detection [52], etc.
+Despite the huge progress [18, 46, 54] following SAM, applying SAM for object detection in crowded scenes is seldom studied.
+In this paper, we investigate the potential of SAM in such cases with two motivations.
+First, SAM is pre-trained on a very large dataset i.e.
+SA-1B that contains most of the common objects and it is thus reasonable to utilize the knowledge to facilitate labeling massive data and training a brand-new detector.
+Second, SAM demonstrates a robust segmentation ability in handling complicated scenes characterized by clustered objects that are difficult for an object detector trained from scratch.
+To this end, we propose Crowd-SAM, a smart annotator powered by SAM for object detection in crowded scenes.
 As depicted in Fig.
@@
 Our method employs dense grids equipped with an Efficient Prompt Sampler (EPS) to cover as many objects as possible at a moderate cost.
-To distinguish the masks from multiple outputs precisely in occluded scenes, we propose a mask selection module, termed Part-Whole Discrimination Network (PWD-Net) that learns to distinguish the output with the highest quality.
-With a lightweight design and fast training schedule, it achieves considerable performance on benchmarks including CrowdHuman[35] and CityPersons[55].
+To distinguish the masks from multiple outputs precisely in occluded scenes, we design a mask selection module, termed Part-Whole Discrimination Network (PWD-Net) that learns to differentiate the output with the highest quality in Intersection over Union (IoU) score.
+With a lightweight model design and fast training schedule, it delivers considerable performance on public benchmarks including CrowdHuman [37] and CityPersons [58].
 Our contributions can be summarized as follows:
   1. 1.
-    We introduce Crowd-SAM - a self-prompted segmentation method for labeling crowded datasets that can produce precise and accurate results with only a few examples.
+    We propose Crowd-SAM, a self-prompted segmentation method, for labeling images containing clustered objects, producing accurate results with only a few examples.
   2. 2.
-    We design two novel components of Crowd-SAM, i.e. EPS and PWD-Net which can effectively unleash SAM’s ability on crowded scenes.
+    We design two novel components of Crowd-SAM, i.e. EPS and PWD-Net, which effectively unleash the ability of SAM on crowded scenes.
   3. 3.
@@
 Object Detection.
-General object detection aims to identify and locate objects and is mainly divided into two categories: one-stage detectors and two-stage detectors.
-One-stage detectors predict bounding boxes and class scores directly from image features[33, 25, 22], while two-stage detectors first generate region proposals and then classify and refine these proposals [34, 10, 9].
-Recently, end-to-end object detectors such as DETR[2, 61, 52] deprecate the hand-crafted modules such as non-maximum suppression (NMS) by adopting one-to-one matching in the training phase, showing great potential in a wide variety of areas.
-However, applying these detectors directly to pedestrian detection tasks will lead to performance degradation due to the fact that pedestrians are often crowded and occluded.
-Early work[30] proposes integrating extra features into a pedestrian detector to explore the low-level visual cues, while follow-up methods [5, 53] attempt to utilize the head areas for better representation learning.
-In [53], an anchor is associated with two targets, the whole body and the head part, to achieve a more robust detector from joint training.
+General object detection aims to identify and locate objects and is mainly divided into two categories: i.e.
+one-stage detectors and two-stage detectors.
+One-stage detectors predict bounding boxes and class scores by using image features[35, 27, 23], while two-stage detectors first generate region proposals and then classify and refine them [36, 10, 9].
+Recently, end-to-end object detectors e.g.
+DETR[2, 63, 55] have replaced the hand-crafted modules such as Non-Maximum Suppression (NMS) by adopting one-to-one matching in the training phase, showing great potential in a wide variety of areas.
+However, applying these detectors directly to pedestrian detection tends to incur performance degradation due to the fact that pedestrians are often crowded with occlusions appearing.
+Early work [32] proposes to integrate extra features into a pedestrian detector to explore low-level visual cues, while follow-up methods [5, 56] attempt to utilize the head areas for better representation learning.
+In [56], an anchor is associated with two targets, the whole body, and the head part, to achieve a more robust detector from joint training.
+AdaptiveNMS [25] adjusts the NMS threshold by predicting the density of pedestrians.
 Alternative methods focus on the design of loss functions to improve the training process.
-For example, RepLoss[43] encourages the prediction consistency of the same target while repels ones from different targets.
+For example, RepLoss [45] encourages the prediction consistency of the same target while repels the ones from different targets.
 Recently, Zheng et al.
-[60] model the relation of queries to improve DETR-based detectors in crowded scenes and achieves remarkable success.
+[62] models the relation of queries to improve DETR-based detectors in crowded scenes and achieves remarkable success.
 Although these works have pushed the boundaries of object detection in crowded scenes to a new stage, they all rely on a large number of labeled samples for training, which is labor-intensive.
@@
 Few-Shot Object Detection (FSOD).
-The task aims to detect objects of novel classes from limited samples.
-FSOD methods can be generally categorized into meta-learning based [48, 15] and fine-tuning based methods[42, 37, 32].
+This task aims to detect objects of novel classes with limited training samples.
+FSOD methods can be roughly classified into meta-learning based [51, 16] and fine-tuning based ones [44, 39, 34].
 Meta-RCNN [51] processes the query and support images in parallel via a siamese network.
-The query’s region of interest (ROI) features are fused with the class prototypes to effectively transfer knowledge learned from the support set.
-TFA[42] proposes a simple two-stage fine-tuning method that only fine-tunes the detector’s last layers.
-FSCE[37] introduces a supervised contrastive loss in the fine-tune stage to mitigate misclassification issues.
-De-FRCN[32] stops the gradient from the RPN and scales the gradient from the R-CNN, followed by a Prototypical Calibration Block to refine the classification scores.
+The Region of Interest (RoI) features of the query are fused with the class prototypes to effectively transfer knowledge learned from the support set.
+TFA [44] proposes a simple two-stage fine-tuning method that only fine-tunes the last layers of the network.
+FSCE [39] introduces a supervised contrastive loss in the fine-tuning stage to mitigate misclassification issues.
+De-FRCN [34] stops the gradient from the RPN and scales the gradient from R-CNN[36], followed by a prototypical calibration block to refine the classification scores.
 Segment Anything Models.
-SAM[19], proposed as a visual foundation model for the segmentation task, is trained on the SA-1B dataset using a semi-supervised learning paradigm.
+SAM [20], a visual foundation model for segmentation tasks, is trained on the SA-1B dataset using a semi-supervised learning paradigm.
 Its exposure to this vast repository of training samples renders it a highly capable class-agnostic model, effectively handling a wide range of objects in the world.
@@
 RSprompt [4] enables SAM to generate semantically distinct segmentation results for remote sensing images by generating appropriate prompts.
-Med-SA[45] proposes a Space-Depth Transpose method to adapt 2D SAM to 3D medical images and a Hyper-Prompting Adapter to achieve prompt-conditioned adaptation.
-Unfortunately, these approaches necessitate a considerable amount of labeled data for effective adaptation, rendering them impractical for crowded scenes where annotation costs are prohibitive.
-Different from them, Per-SAM[54] and Matcher [26] teaches SAM to recognize specified objects with only one or few instances by extracting training-free similarity priors.
-SAPNet [44] proposes a weakly-supervised pipeline for instance segmentation.
-Although these approaches reduce the data requirement, they still lag behind the demands of crowded scenes, such as pedestrian detection, particularly in occlusion scenes.
+Med-SA [47] presents a space-depth transpose method to adapt 2D SAM to 3D medical images and a hyper-prompting adapter to achieve prompt-conditioned adaptation.
+Unfortunately, these approaches necessitate a considerable amount of labeled data for effective adaptation, making them impractical for crowded scenes where annotation costs are prohibitive.
+Different from them, Per-SAM [57] and Matcher [28] teach SAM to recognize specified objects with only one or few instances by extracting training-free similarity priors.
+SAPNet [46] builds a weakly-supervised pipeline for instance segmentation.
+Although these approaches reduce data requirements, they still lag behind the demands of crowded scenes, such as pedestrian detection, particularly with occlusions.
 Figure 2: The pipeline of Crowd-SAM shows the interaction between different modules.
@@
 For simplicity, the projection adapter of DINO is dismissed.
-3 Our Approach
+3 Method
 3.1 Preliminaries
-SAM [19] is a promising segmentation model that comprises three main components: (a) an image encoder responsible for feature extraction; (b) a prompt encoder designed for encoding the geometric prompts provided by the users; and (c) a lightweight mask decoder that predicts the masks conditioned on the given prompts.
+SAM [20] is a powerful and promising segmentation model that comprises three main components: (a) an image encoder responsible for feature extraction; (b) a prompt encoder designed to encode the geometric prompts provided by users; and (c) a lightweight mask decoder that predicts the masks conditioned on the given prompts.
 Leveraging extensive training data, SAM demonstrates impressive zero-shot segmentation performance across various benchmarks.
-In particular, SAM utilizes points and boxes as prompts to specify the interested regions.
-DINO [3] represents a family of self-supervised vision transformers [7] designed for general-purpose applications.
-During its training, DINO employs a self-distillation approach akin to BYOL [11], fostering the learning of robust feature representations.
-DINOv2 [31] enhances the foundation of DINO by integrating additional pre-training tasks, improving its scalability and stability, especially in large models like ViT-H with 1 billion parameters.
-Thanks to its enhancement, DINOv2 demonstrates strong feature representation ability, especially for the task of semantic segmentation.
-3.2 Problem Definition and Framework
+In particular, SAM makes use of points and boxes as prompts to specify interested regions.
+DINO [3] represents a family of Vision Transformers (ViT) [7] learned in a self-supervised manner designed for general-purpose applications.
+During its training, DINO employs a self-distillation strategy akin to BYOL [11], fostering the learning of robust representations.
+DINOv2 [33] strengthens the foundation of DINO by integrating several additional pre-training tasks, improving its scalability and stability, especially for large models e.g.
+ViT-H (1 billion parameters).
+Thanks to its enhancement, DINOv2 shows a strong representation ability, in particular for the task of semantic segmentation.
+3.2 Problem Definition and Overall Framework
+Problem Definition.
 As shown in Fig.
-1, our goal is to detect pedestrians in crowded scenes with few annotated data.
+1, our goal is to detect objects (e.g.
+pedestrians) in crowded scenes with few annotated data.
 We formulate this problem as a one-class few-shot detection task.
-A common few-shot approach is to split the data into the base split and the novel split.
-Differently, we directly use the data of the target class to train our model, since the foundation models have already been trained on massive data.
-In particular, we use segmentation masks as intermediate results which can be easily converted to bounding boxes.
-During the training and evaluation process, only box annotations are provided.
-Naive study on SAM auto generator.
-Before we introduce our method, we take a deeper look into how the prompt number affects SAM’s performance in crowded scenes.
-To this end, we conduct several naive studies on CrowdHuman[35] with SAM’s auto generator, which utilizes grid points to search every region.
+A common few-shot pipeline is to divide data into the base split and the novel split.
+Differently, we directly use the data of the target class for model training, as the foundation models have already been trained on massive data.
+In particular, we employ segmentation masks as intermediate results which can be easily converted to bounding boxes.
+During the training and evaluation processes, only box annotations are provided.
+Naive Study on SAM Auto-generator.
+The prompt number affects the performance of SAM and we analyze this issue for crowded scenes.
+In this case, we conduct several naive studies on CrowdHuman [37] with the auto generator of SAM, which utilizes grid points to search every region.
 Tab.
-1 conveys three key observations: (1) dense grids are necessary for crowded scenes; (2) the ambiguity of point prompts and class-agnostic prompts distribution incur many false positives (FPs); (3) decoding time is a non-negligible burden when grid size is large.
-It is concluded that the dense prompts and FP removal are key aspects in designing SAM-based methods for detection/segmentation tasks in crowded scenes.
-N 16 32 64 128 Oracle
+1 conveys three key observations: (1) dense grids are necessary for crowded scenes; (2) the ambiguity of distributions of point prompts and class-agnostic prompts incur many false positives (FPs); (3) the decoding time is a non-negligible burden when the grid size is large.
+Therefore, the dense prompts and FP removal are key aspects in designing SAM-based methods for detection/segmentation tasks in crowded scenes.
+Table 1: Comparison in terms of recall, average FPs, and decoding time ($T$) of different grid sizes ($N_{G}$) adopted by the SAM generator on CrowdHuman [37].
+The oracle model derives the prompts by computing the center of ground truth boxes.
+The decoding time is collected on a 3090 Ti GPU card.
+$N_{G}$ 16 32 64 128 Oracle
 Recall 33.6 58.0 63.4 76.0 91.4
-avg. FP 51 112 227 485 -
+avg. FPs 51 112 227 485 -
 $T$ (s) 0.059 0.22 0.83 3.2 0.045
-Table 1: Comparisons of different grid sizes adopted by SAM generator on Crowd-Human[35].
-Oracle model derives the prompts by computing the center of the ground truth boxes.
-T is the decoding time collected on a 3090 Ti GPU card.
-Framework.
-Inspired by the above studies, we equip SAM [19] with several proper components to achieve an accurate and efficient annotation framework, as illustrated in  Fig.
+Overall Framework.
+Inspired by the studies above, we equip SAM [20] with several proper components to achieve an accurate and efficient annotation framework, as illustrated in Fig.
 2.
-To accurately locate pedestrians, we employ the foundational model DINOv2[31] to predict a semantic heat map, a task that can be formulated with a simple binary classifier.
-To discriminate the output masks that is a mixture of correct masks, backgrounds, and part-level masks, we adopt Part-Whole Discrimination Network that takes as input both the learned tokens from SAM and the semantic-rich tokens from DINOv2 to re-evaluate all the outputs.
+To accurately locate clustered objects, we employ the foundational model DINOv2[33] to predict a semantic heat map, a task that can be formulated with a simple binary classifier.
+To discriminate the output masks that is a mixture of correct masks, backgrounds, and part-level masks, we design Part-Whole Discrimination Network (PWD-Net) that takes as input both the learned tokens from SAM and the semantic-rich tokens from DINOv2 to re-evaluate all the outputs.
 Finally, to handle the redundancy brought by the use of dense grids, we propose an Efficient Prompt Sampler (EPS) to decode the masks at a moderate cost.
 We introduce the details of our methods in the following sections.
-3.3 Class-specific Prompts Generation
-Generating a unique point prompt for each pedestrian in the dense scenes is de facto a non-trivial problem.
-Thus, we take a step back and study how to detect pedestrians with multiple prompts associated with one person, and then use the proper post-processing methods to remove duplicates.
-To this end, we adopt a heat-map-based prompt generation pipeline that classifies the regions first and generates prompts from the positive regions.
-For the input image $I\in\mathbb{R}^{H\times W\times 3}$, we use a pre-trained image encoder $E_{D}$ to extract semantic rich features first.
-To better transfer the pre-trained features to pedestrian segmentation, we add an MLP block after the final output layer and thus get the adapted features $F_{DINO}\in\mathbb{R}^{{H\over s}\times{W\over s}\times C}$, where $s$ is the patch size of DINOv2 [31] and $C$ is the output channel.
-Then, we use a segmentation head $\text{Head}_{CLS}$ to classify $F_{DINO}$ pixel by pixel, resulting in a heatmap ^$\hat{\mathbb{H}}\in[0,1]^{{H\over s}\times{W\over s}}$ that indicates the locations of pedestrians as ^$\hat{\mathbb{H}}=\text{Head}_{CLS}(F_{DINO})$.
-Given only the bounding box annotations $B\in[0,1]^{N\times 4}$, where $N$ is the number of targets, this binary segmentation head can be optimized with box-level supervision.
-However, the coarse boundaries would incur considerable points scattering in the background regions.
-To alleviate this issue, we use SAM [19] to obtain high-quality mask-level pseudo labels, which is illustrated in  Fig.
+3.3 Class-specific Prompt Generation
+Generating a unique point prompt for each object (e.g.
+pedestrian) in crowded scenes is de facto a non-trivial problem.
+Thus, we take a step back and study how to detect objects with multiple prompts associated with one object and apply the proper post-processing techniques to remove duplicates.
+To this end, we adopt a heatmap-based prompt generation pipeline that initially classifies the regions and then generates prompts from the positive regions.
+For the input image $I\in\mathbb{R}^{H\times W\times 3}$, we first use a pre-trained image encoder $E_{D}$ to extract semantic rich features.
+To better transfer the pre-trained features to object segmentation, we add an MLP block after the final output layer and thus obtain the adapted features $F_{DINO}\in\mathbb{R}^{{H\over s}\times{W\over s}\times C}$, where $s$ is the patch size of DINOv2 [33] and $C$ is the output channel.
+Then, we employ a segmentation head $\text{Head}_{CLS}$ to classify $F_{DINO}$ pixel by pixel, resulting in a heatmap ^$\hat{\mathbb{H}}\in[0,1]^{{H\over s}\times{W\over s}}$ that indicates the locations of objects as ^$\hat{\mathbb{H}}=\text{Head}_{CLS}(F_{DINO})$.
+Given only the bounding box annotations $B\in[0,1]^{N_{B}\times 4}$, where $N_{B}$ is the number of targets, this binary segmentation head can be optimized with box-level supervision.
+However, the coarse boundaries tend to incur considerable points scattering in background regions.
+To alleviate this issue, we use SAM [20] to generate high-quality mask-level pseudo labels, which is illustrated in Fig.
 4, with ground truth (GT).
 The decoded masks are then merged into a single foreground mask $\mathbb{H}\in\{0,1\}^{256\times 256}$.
-We use dice loss for training the adapter and segmentation head with the generated pseudo masks as follows:
+We use the dice loss for training the adapter and segmentation head with the generated pseudo masks as follows:
 $^{\cal L}_{fg}=dice(f(\hat{\mathbb{H}}),\mathbb{H}),$ (1)
-where $f$ is a up-sampling function that resizes $\mathbb{H}$ to $256\times 256$.
-During inference, we use a threshold $t$ for mask binarization, which is simply set 0.5 in our experiments.
+where $f$ is an up-sampling function that resizes $\mathbb{H}$ to $256\times 256$.
+During inference, we add a threshold $t$ for mask binarization, which is simply set at 0.5 in our experiments.
 The binarized masks are mapped to point prompts $P_{G}$ which only contain those in positive regions.
@@
 Given the proposals generated in Sec.
-3.3, this section aims to decode the dense prompts efficiently and discriminate the generated masks accurately.
+3.3, our further aim is to efficiently decode the dense prompts and accurately discriminate the generated masks.
 As depicted in Fig.
-3, each instance contains many prompts due to the density of grids.
-Supposing only one in-position prompt is required for mask prediction for SAM, decoding all the prompts would lead to not only a waste of computation but also cause more FPs due to some badly located prompts.
-Figure 3: Illustration of EPS. The PWD-Net produce valid masks with a confidence threshold $T$. In each iteration, we prune prompts that fall inside valid masks (with a cross above).
-Efficient Prompt Sampler.
-To address this challenge, we introduce the Efficient Prompt Sampler (EPS), which dynamically prunes prompts based on the confidence of decoded masks.
+3, each instance contains a set of prompts due to the density of grids.
+Supposing that only one in-position prompt is required for mask prediction in SAM, decoding all the prompts would lead to not only a waste of computation but also more FPs for some poorly located prompts.
+Figure 3: Illustration of EPS. PWD-Net produces valid masks with a threshold. In each iteration, we prune prompts (with a cross above) that fall inside valid masks .
+Efficient Prompt Sampler (EPS).
+To address this challenge, we introduce EPS, which dynamically prunes prompts based on the confidence of decoded masks.
 This method is elaborated in Algorithm 1.
@@
 The sampled prompts $P_{B}$ are then appended to the output points list $P_{S}$.
-Subsequently, we employ SAM Generator with the batched prompts $P_{B}$ to produce masks $M$.
+Subsequently, we employ the SAM generator with the batched prompts $P_{B}$ to produce masks $M$.
 Furthermore, we aggregate discriminant confidence scores $S$ from PWD-Net, which are employed to select valid masks $M^{\prime}$ using a score threshold $T$, by using $M^{\prime}=M[S>T]$.
-A detailed exposition of PWD-Net will be provided in the subsequent section.
-The valid masks represent regions that are believed to be well segmented, thus we remove the points that are already covered by any mask $m$ in $M^{\prime}$ from the prompts list $P_{G}$.
+Refer to a detailed exposition of PWD-Net in the subsequent section.
+The valid masks represent regions that are believed to be well segmented, and we thus remove the points that have already been covered by any mask $m$ in $M^{\prime}$ from the prompts list $P_{G}$.
 The iteration halts when $P_{G}$ is empty.
-Furthermore, EPS establishes a stopping criterion with a parameter $K$, terminating the sampling process once the total sample count reaches this limit.
+Additionally, EPS establishes a stopping criterion with a parameter $K$, terminating the sampling process once the total sample count reaches this limit.
 This parameter is instrumental in managing the overall decoding cost.
-Input: Feature extracted from image $I$ by image encoder of SAM: $F_{SAM}$; Generated prompt list: $P_{G}=\{p_{1},p_{2},...,p_{N}\}$; SAM Generator: $G$; Mask confidence threshold: $T$; Target sampled prompt list size: $K$;
-Output: Sampled prompt list: $P_{S}$; Valid masks list: $M_{S}$.
+In our experiments, $K$ is empirically set to 500 to strike a balance between efficiency and recall.
+Input: feature extracted from image $I$ by image encoder of SAM: $F_{SAM}$; generated prompt list: $P_{G}=\{p_{1},p_{2},...,p_{N}\}$; SAM generator: $G$; mask confidence threshold: $T$; target sampled prompt list size: $K$
+Output: sampled prompt list: $P_{S}$; valid mask list: $M_{S}$
 1 $P_{S}\leftarrow\emptyset$;
@@
 Algorithm 1 Process of Efficient Prompt Sampler (EPS).
-Part-Whole Discrimination Network.
-Given the raw masks predicted by SAM conditioned on the sampled batch of prompts $P_{B}$, we are inspired to design an automatic selector for choosing the best fitting mask.
-It should refine the outputs in two aspects: (i).
-help to refine the output IoU score according to the quality of related masks, if they are positive samples; (ii).
-help suppress the scores of samples that fall in the background regions.
-usion of those associated with the background.
+Part-Whole Discrimination Network (PWD-Net).
+Given the raw masks predicted by SAM conditioned on the sampled batch of prompts $P_{B}$, we design an automatic selector for choosing the best-fitting mask.
+It is expected to optimize the outputs in two aspects: (i) refining the output IoU score according to the quality of related masks if they are positive samples and (ii) suppressing the scores of samples that fall in background regions.
 Illustrated in Fig.
-2, for masks generated corresponding to $N$ prompts, we leverage the Mask Tokens and IoU Tokens within the mask decoder of SAM along with the sophisticated features extracted by the self-supervised pre-trained model DINOv2 [31].
-${\cal M}$ and ${\cal U}$ are responsible for mask decoding and IoU prediction in SAM Mask Decoder, respectively.
-Thus, we suppose that they contain shape-aware information, which is helpful in discriminating the mask.These components enable us to compute a discriminant confidence score $S$ for each specific prompt in a few-shot approach.
+2, for the masks generated corresponding to $N$ prompts, we leverage the Mask Tokens and IoU Tokens within the mask decoder of SAM along with the sophisticated features extracted by the self-supervised pre-trained model DINOv2 [33].
+${\cal M}$ and ${\cal U}$ are responsible for mask decoding and IoU prediction in the SAM Mask Decoder, respectively.
+Thus, we suppose that they contain shape-aware information, which is helpful in discriminating the mask.
+These components enable us to compute a discriminant confidence score $S$ for each specific prompt in a few-shot scenario.
 Initially, the refined IoU score $S_{iou}$ is computed as follows:
 $S_{iou}=\text{Head}_{par}(\text{Concat}(\text{Repeat}({\cal U}),{\cal M}))+\text{Head}_{IoU}({\cal U}),$ (2)
-where $\text{Head}_{par}$ is a parallel IoU head, consisting of a MLP block, ${\cal U}\in\mathbb{R}^{N\times 1\times C}$ and ${\cal M}\in\mathbb{R}^{N\times 4\times C}$ denote the IoU Tokens and Mask Tokens respectively.
+where $\text{Head}_{par}$ is a parallel IoU head, consisting of an MLP block; ${\cal U}\in\mathbb{R}^{N\times 1\times C}$ and ${\cal M}\in\mathbb{R}^{N\times 4\times C}$ denote the IoU Tokens and Mask Tokens respectively.
 Notice in Eq.
-2, the $S_{iou}$ is a sum of outputs from two individual heads, the parallel adapter head $\text{Head}_{par}$ and the original $\text{Head}_{IoU}$.
+2, $S_{iou}$ is a sum of outputs from two individual heads, the parallel adapter head $\text{Head}_{par}$ and the original $\text{Head}_{IoU}$.
 We freeze the parameters of $\text{Head}_{IoU}$ to avoid overfitting in few-shot learning.
-Moreover, since these two tokens, ${\cal M}$ and ${\cal U}$ have different shapes, we repeat ${\cal U}$ by 4 times and then concat these two tokens in the channel dimension.
+Moreover, since these two tokens, ${\cal M}$ and ${\cal U}$ have different shapes, we repeat ${\cal U}$ by 4 times and concatenate these two tokens in the channel dimension.
 The refined score $S_{iou}\in\mathbb{R}^{N\times 4}$ encapsulates the quality of the generated masks by assessing the texture-aware feature from ${\cal M}$ and ${\cal U}$.
-Secondly, by harnessing the semantic feature embedded within the self-supervised pre-trained model DINOv2, along with the mask data $\hat{M}$, we calculate the discriminant score $S_{cls}$ :
+Further, by harnessing the semantic feature embedded in the self-supervised pre-trained model DINOv2, along with the mask data $\hat{M}$, we calculate the discriminant score $S_{cls}$ :
 $S_{cls}=\sigma({\text{Head}_{CLS}({\cal O})}),{\cal O}=\text{Pool}(d(\text{Softmax}(\hat{M}))\circ F_{DINO}).$ (3)
-Here, we use ${\cal O}\in\mathbb{R}^{N\times C}$ to denote the extracted Semantic Token, $\hat{M}\in\mathbb{R}^{N\times H\times W\times 4}$ and $F_{DINO}\in\mathbb{R}^{N\times h\times w\times C}$ to represent the masks generated by SAM[29] and the features extracted by the DINOv2[31] model, respectively.
-We denote $d$ as a down-scale function that resizes the mask to $h\times w$ to be consistent with $F_{DINO}$.
-Pool is a global pooling function that conducts mean pooling on the x-axis and y-axis.The discriminant score $S_{cls}\in\mathbb{R}^{N\times 4}$ predicts whether masks belong to the foreground or background.
-The $\text{Head}_{CLS}$ shares the same parameters with the binary classifier introduced in  Sec.
+Here, ${\cal O}\in\mathbb{R}^{N\times C}$ denotes the extracted Semantic Token, and $\hat{M}\in\mathbb{R}^{N\times H\times W\times 4}$ and $F_{DINO}\in\mathbb{R}^{N\times h\times w\times C}$ represent the masks generated by SAM[31] and the features extracted by DINOv2, respectively.
+We denote $d$ as a down-scale function that resizes the mask to $h\times w$, consistent with $F_{DINO}$.
+Pool is a global pooling function that conducts mean pooling on the x-axis and y-axis and $\circ$ is the Hadamard product.
+The discriminant score $S_{cls}\in\mathbb{R}^{N\times 4}$ predicts whether masks belong to the foreground or background.
+$\text{Head}_{CLS}$ shares the same parameters with the binary classifier introduced in Sec.
 3.3.
-Finally, we calculate a joint score of discrimination and estimate quality for masks by simply multiplying the two scores together: $S=S_{iou}\cdot S_{cls}$.
-During training, prompts sampled within real masks are utilized as input.
-Regarding prompts within the foreground, the discriminant confidence score should aim to accurately predict the Intersection over Union (IoU) of generated and real masks.
-Conversely, for prompts within the background, the score should ideally output 00, indicating their association with the background.
+Finally, we calculate a joint score of discrimination and estimate the quality for masks by simply multiplying the two scores: $S=S_{iou}\cdot S_{cls}$.
+During training, prompts sampled from real masks are taken as input.
+Regarding the prompts within the foreground, the discriminant confidence score aims to accurately predict the IoU of the generated and real masks.
+Conversely, for the prompts within the background, the score is ideally 00.
 Hence, the loss function for this aspect is formulated as follows:
-$s^{i}_{target}=\begin{cases}mIoU(m^{i},m^{i}_{GT}),&m^{i}_{GT}\in M^{bg}_{GT},\\ 0,&m^{i}_{GT}\in M^{fg}_{GT},\end{cases}$ (4)
+$s^{i}_{target}=\begin{cases}IoU(m^{i},m^{i}_{GT}),&m^{i}_{GT}\in M^{bg}_{GT},\\ 0,&m^{i}_{GT}\in M^{fg}_{GT},\end{cases}$ (4)
 ${\cal L}_{iou}=\text{MSE}(S,S_{target})$ (5)
@@
 3.5 Training and Inference
-The total training loss of our framework is the combination of  Eq. 1 and Eq. 5:
+The total training loss of the entire framework combines Eq. 1 and Eq. 5:
 ${\cal L}={\cal L}_{fg}+{\cal L}_{iou}$ (6)
-During inference, we select the mask with the highest confidence score $S$ among the $4$ masks as the output of PWD-Net, which we denote as $M\in\mathbb{R}^{N\times H\times W}$ and $S_{o}\in\mathbb{R}^{N}$.
-We adopt a window cropping strategy for better performance on small objects as [19].
-This strategy slices the whole image into overlapping crops where each crop is processed individually.
-The final results are merged from the outputs of each crop.
-We apply non-maximum suppression on the final outputs to remove duplicate proposals.
+At inference, we select the mask with the highest confidence score $S$ among the $4$ masks as the output of PWD-Net, which we denote as $M\in\mathbb{R}^{N\times H\times W}$ and $S_{o}\in\mathbb{R}^{N}$.
+We adopt a window cropping strategy to enhance the performance on small objects as [20].
+This strategy slices the whole image into overlapping crops where each crop is individually processed.
+The final results are merged from the outputs of each crop and here we apply NMS to remove duplicate proposals.
 4 Experiments
+Table 2: Comparative results (%) on CrowdHuman [37] val.
+All the SAM-based methods adopt ViT-L [7] as the pre-trained backbone (SRCNN denotes Sparse R-CNN [40], a baseline in [62] and * represents using the multi-crop trick).
 Method Backbone #Shot AP MR-2 Recall Secs/Img
@@
 Crowd-SAM* ViT-L[7] 10 78.4 74.8 85.6 8.1
-Table 2: Results on CrowdHuman[35] val set.
-All SAM-based methods adopt ViT-L[7] as the pre-trained backbone.
-SRCNN represents Sparse R-CNN[38] which is a baseline to [60].
-* represents using the multi-crop trick.
 Datasets.
@@
 It contains 15,000, 4,370, and 5,000 images for training, validation, and testing, respectively.
-We also test our method on CityPersons[55] for a realistic urban-scene evaluation.
-Additionally, to further investigate our method, we utilize OCC-Human [58] for evaluation, which is specially designed for occluded persons.
-For these pedestrian datasets, we utilize visible annotations (only including visible areas of an object) for training and evaluation.
-To validate the extensibility of Crowd-SAM, we also devise a multi-class version of Crowd-SAM by adding a multi-class classifier.
-We utilize 0.1 % percent of the COCO [23] trainval set for training and COCO val set for validation.
-We also validate our methods on an occluded subset of COCO, COCO-OCC [16] that is extracted by selecting images whose objects have a high overlapping ratio.
+We also evaluate our method on CityPersons[58] for a realistic urban-scene scenario.
+Additionally, we utilize OCC-Human [61], which is specially reputed for occluded persons.
+For these pedestrian datasets, we use visible annotations (only including visible areas of an object) for training and evaluation.
+To validate the extensibility of Crowd-SAM, we further devise a multi-class version of Crowd-SAM by adding a multi-class classifier.
+We employ 0.1 % percent of the COCO [24] trainval set for training and the COCO val set for validation.
+Besides, we validate our method on a subset with occluded objects on COCO, i.e.
+COCO-OCC [17], extracted by selecting the images whose objects have a high overlapping ratio.
 Implementation Details.
 We utilize SAM (ViT-L) [20] and DINOv2 (ViT-L) [33] as the base models for all experiments.
-In the fine-tuning stage, their parameters are all frozen to avoid over-fitting.
-Instead of real GT, we use pseudo masks generated by SAM (ViT-L) as $S_{target}$ to supervise the learning of PWD-Net in Eq.
+In the fine-tuning stage, all their parameters are frozen to avoid over-fitting.
+Instead of real GT, we use the pseudo masks generated by SAM as $S_{target}$ to supervise the learning of PWD-Net in Eq.
 5.
@@
 4(c).
-We randomly pick points from the pseudo masks as positive training samples and points from the background part as negative training samples.
+We randomly pick the points from the pseudo masks as positive training samples and the ones from the background as negative training samples.
 In the training process, we use Adam[19] with a learning rate of $10^{-5}$, a weight decay of $10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.99$ for optimization.
-We train our module for 2000 iterations with a batch size of 1, which can be done on a single GTX 3090 Ti GPU in several minutes.
+We train our module for 2,000 iterations with a batch size of 1, which can be done on a single GTX 3090 Ti GPU in several minutes.
 For more details, please refer to the Appendix.
-Evaluation metrics. Following [60], we use AP with IoU threshold 0.5, MR-2, and Recall for our evaluation metrics. Generally, a higher AP, Recall, and lower MR-2 indicates a better performance.
+Evaluation Metrics. Following [62], we use AP with IoU threshold at 0.5, MR-2, and Recall as metrics. Generally, a higher AP, Recall, and lower MR-2 value indicates better performance.
 4.1 Experimental Results on Pedestrian Detection
-For a fair comparison, we re-implement the involved methods[56, 41, 60, 52, 32] with a 2×2$\times$2 × schedule using the visible annotations in the CrowdHuman [35] and CityPersons [55] dataset.
+For fair comparison, we re-implement the counterparts[59, 43, 62, 55, 34] with a 2×2$\times$2 × schedule using visible annotations in the CrowdHuman [37] and CityPersons [58] datasets.
 Main Results.
-We compare Crowd-SAM with most related methods including fully supervised object detectors[41, 56, 52, 60], few shot object detectors[32] and SAM-based approaches[26, 19] for comprehensive understanding.
-Notice we use the visible annotation for which derives different results from those on full box annotation.
-As shown in  Tab.
-2, with only 10 labeled images, our method achieves comparable performance with full-supervised object detectors whose best result is 86.7% AP, achieved by DINO[52].
+We compare Crowd-SAM with most related methods including fully-supervised object detectors [43, 59, 55, 62], few shot object detectors [34] and SAM-based methods [28, 20].
+Notice that we use visible annotations for which derive different results from those on full box annotations.
+As Tab.
+2 shows, with only 10 labeled images, Crowd-SAM achieves comparable performance with full-supervised object detectors whose best result is 86.7% AP, delivered by DINO[55].
 Particularly, Crowd-SAM outperforms an advanced anchor-free detector FCOS[43] by 2.1% AP.
-These results indicate that with proper adaptation techniques, SAM can achieve very competitive performance on complex pedestrian detection datasets like CrowdHuman.
-On the other hand, Crowd-SAM achieves SOTA performance on few-shot detection settings and outperforms all few-shot object detectors with a significant margin.
-DeFRCN [32] is a well-established few-shot object detector equipped with ResNet-101[13] that achieves 46.4% in AP.
-Notably, our method outperforms it by 32%, indicating the superiority of Crowd-SAM in the few-shot detection setting.
-The qualitative comparisons between ours and De-FRCN are depicted in  Fig.
+These results indicate that by using proper adaptation techniques, SAM can reach very competitive performance on complex pedestrian detection datasets like CrowdHuman.
+On the other side, Crowd-SAM achieves SOTA performance on few-shot detection settings and outperforms all the few-shot object detectors by a significant margin.
+DeFRCN [34] is a well-established few-shot object detector equipped with ResNet-101 [14] that reports 46.4% in AP.
+Notably, our method exceeds it by 32%, indicating the superiority in the few-shot detection setting.
+The qualitative comparison between Crowd-SAM and De-FRCN is depicted in Fig.
 4.
-For SAM-based approaches, our method largely leads the SAM baseline by 6% (with multi-crop) and 18.3% (w/o multi-crop) in the recall.
-Crowd-SAM is also superior to other methods such as Matcher[26].
+For SAM-based methods, our Crowd-SAM largely leads the SAM baseline by 6% (with multi-crop) and 18.3% (w/o multi-crop) in Recall.
+Besides, Crowd-SAM is also superior to the other methods such as Matcher [28].
 Results on OccHuman and CityPersons.
-To investigate the performance of Crowd-SAM in occluded scenes, we compare it with an advanced supervised method Pose2Seg [58], and the results are shown in Tab.
+To investigate the performance of Crowd-SAM in occluded scenes, we compare it with an advanced counterpart Pose2Seg [61], and the results are shown in Tab.
 3.
-It is noteworthy that Pose2Seg is a full supervised detector while Crowd-SAM is a few-shot detector.
+It is noteworthy that Pose2Seg is a fully-supervised detector while Crowd-SAM is a few-shot detector.
 As can be seen from the results, Crowd-SAM leads Pose2Seg by 9.2% AP which demonstrates its robustness in occluded scenes.
-Also, we apply our method to a not-so-crowded but more realistic urban dataset, CityPersons [55], as reported in Tab.
+Also, we apply our method to a not-so-crowded but more realistic urban dataset, i.e.
+CityPersons [58], as reported in Tab.
 4.
@@
 In conclusion, our method remains competitive compared to advanced few-shot object detectors, even though it is not specifically designed for sparse scenes.
-These results illustrate that our method has fully unleashed the power of vision foundation model models, i.e. SAM and DINO, in all of the crowded, occluded, and urban scenes.
+These results illustrate that Crowd-SAM well unleashes the power of vision foundation models, i.e. SAM and DINO, in all of the crowded, occluded, and urban scenes.
+Table 3: Comparative results (%) on OccHuman val, where APM and APH represent AP in moderate and hard cases according to occlusion ratios, respectively.
 Method Backbone AP APM APH
-Mask R-CNN [12] R50-FPN 16.3 19.4 11.3
-Pose2Seg [58] R50-FPN 22.2 26.1 15.0
-Crowd-SAM ViT-L 31.4 26.5 17.7
-Table 3: Comparative results on OccHuman [58] val. APM and APH represent AP in moderate and hard cases according to occlusion ratio, respectively.
+Mask R-CNN [13] ResNet50-FPN [14] 16.3 19.4 11.3
+Pose2Seg [61] ResNet50-FPN [14] 22.2 26.1 15.0
+Crowd-SAM ViT-L [7] 31.4 26.5 17.7
+Table 4: Comparative results (%) on CityPersons val.
 Method # Shot AP MR-2
@@
 Crowd-SAM 50 33.3 31.7
-Table 4: Comparative results on CityPersons[55] val.
+Table 5: Comparative results (%) on COCO val and COCO-OCC.
 Methods Backbone COCO-OCC COCO
 AP AP50 AP AP50
-Faster R-CNN [34] R-50-FPN 29.7 50.0 33.5 53.7
-BCNet [16] R-50-FPN 31.7 51.1 34.6 54.4
-Crowd-SAM ViT-L 20.6 31.5 22.0 33.7
-Table 5: Comparative results on COCO [23] and COCO-OCC [16] val.
+Faster R-CNN [36] ResNet50-FPN [14] 29.7 50.0 33.5 53.7
+BCNet [17] ResNet50-FPN [14] 31.7 51.1 34.6 54.4
+Crowd-SAM ViT-L [7] 20.6 31.5 22.0 33.7
 4.2 Experimental Results on Multi-class Object Detection
-To further explore the extensibility of our methods in a more popular setting, we devise a multi-class Crowd-SAM.
-The multi-class version of Crowd-SAM is slightly different by replacing the binary classifier with a multi-class classifier.
-We validate multi-class Crowd-SAM on COCO [23], a widely adopted common object detection benchmark, and COCO-OCC [16] which is a split of COCO that mainly consists of images with high occlusion ratio.
-We compare our method with two supervised detectors, Faster R-CNN [34] and BCNet [16], and report the results in Tab.
+To further explore the extensibility in a more popular setting, we devise a multi-class Crowd-SAM.
+The multi-class version of Crowd-SAM is slightly different by replacing the binary classifier with a multi-class one.
+We then validate multi-class Crowd-SAM on COCO [24], a widely adopted common object detection benchmark, and COCO-OCC [17] which is a split of COCO that mainly consists of images with a high occlusion ratio.
+We compare our method with two supervised detectors, i.e.
+Faster R-CNN [36] and BCNet [17], and report the results in Tab.
 5.
-It is seen that our Crowd-SAM is comparable to supervised detectors in both datasets and drops only 1.4 APwhen comparing COCO and COCO-OCC.
-This minor drop indicates that our method is robust to the occlusion.
+It can be seen that our Crowd-SAM is comparable to the supervised detectors on both datasets and drops only 1.4 APwhen comparing those of COCO and COCO-OCC.
+This minor drop indicates that our method is robust to occlusions.
+Table 6: Ablation results (%) on the main components in Crowd-SAM.
+FG location represents the use of the binary classifier on DINOv2 features.
+The last line represents a SAM baseline with a standard $32\times 32$ grid as inputs.
 FG loc. EPS PWD-Net Multi-Crop AP MR-2 Recall
@@
 8.7 99.8 70.1
-Table 6: Ablation studies on the main components in Crowd-SAM.
-FG location represents the use of binary classifier on DINOv2 features.
-The last line represents a SAM baseline with a standard $32\times 32$ grid as inputs.
 4.3 Ablation Studies
-Ablation on the modules.
-We conduct ablation studies on the key components of Crowd-SAM - foreground location, EPS, and PWD-Net - to validate their effectiveness.
+Ablation on Modules.
+We conduct ablation studies on the key components of Crowd-SAM, including foreground location, EPS, and PWD-Net, to validate their effectiveness.
 In Tab.
-6, the performance on AP significantly drops by 7.4% and recall by 8.5% when the FG location is removed, indicating the importance of restricting foreground areas.
+6, the performance in AP significantly drops by 7.4% and recall by 8.5% when FG location is removed, indicating the importance of restricting foreground areas.
 As for EPS, once it is removed, the AP drops by 0.6%.
-We suppose EPS not only accelerates the sampling process of dense prompts but also helps focus on the difficult part of the images, which has ambiguous semantics.
-We also compare EPS with other batch iterators in  Tab.
+We suppose that EPS not only accelerates the sampling process of dense prompts but also helps focus on the difficult part of the image, which conveys ambiguous semantics.
+We also compare EPS with some other batch iterators in Tab.
 7.
-Moreover, we find the PWD-Net is indispensable and when removed, the AP dramatically drops to only 17.0%.
-Finally, we point out that multi-cropping is a strong method to enhance the performance which contributes 7.4 % AP to the final performance.
-Overall, these results prove that all components are essential.
+Besides, we find that PWD-Net is indispensable and when removed, the AP dramatically drops to 17.0%.
+Finally, we point out that multi-cropping is a strong trick to enhance the performance which contributes 7.0% AP to the final performance.
+Overall, these results prove that all the components are essential.
+Table 7: Comparison (%) of different samplers on CrowdHuman [37] val. Full means using all prompts. OOM represents out-of-memory errors which occur when the GPU memory is all consumed.
 Full Random EPS
@@
 256 $\times$ 256 500 OOM 68.9 73.0 72.3 78.0
-Table 7: Comparisons of different sampler. Full means using all prompts. OOM represents out-of-memory error which occurs when the GPU memory is all consumed.
 Ablation on EPS.
 We conduct a comprehensive study on EPS by comparing it with several variants.
-We forward each image only once to avoid the side effects caused by multi-crop.
+We forward each image only once to avoid the extra latency caused by multi-crop.
 We use a default sampler that iterates through all training samples and a random iterator with a halting threshold $K$ as counter-parts.
 As reported in Tab.
-7, AP and Recall increase with the grid size for all three samplers.
-However, the default sampler meets an out-of-memory error when the grid size reaches 128, preventing it from being adopted in this setting.
-As for the random sampler, its performance is constrained by K and a larger grid size than 64 only leads to limited improvement, e.g.
+7, the AP and Recall increase with the grid size for all three samplers.
+However, the default sampler suffers an out-of-memory error when the grid size reaches 128, preventing it from being adopted in this setting.
+As for the random sampler, its performance is constrained by $K$ and a grid size larger than 64 only leads to limited improvement, e.g.
 0.1% AP.
-On the contrary, our EPS can benefit from a much larger grid size and achieves a better performance especially when the grid size is large.
+On the contrary, our EPS benefits from a much larger grid size and achieves a better result.
+Table 8: Ablation results (%) on the design of PWD-Net.
+${\cal M}$, ${\cal U}$, and ${\cal O}$ represent the mask token, IoU token, and semantic token, respectively.
+For the IoU head, $F$ means freezing the original IoU head and training a parallel one, and $T$ indicates tuning the original IoU head.
 ${\cal M}$ ${\cal U}$ ${\cal O}$ IoU head AP MR-2
@@
 $F$ 38.4(-40.0) 95.8(+21.0)
-Table 8: Ablation studies on the design of PWD-Net.
-${\cal M}$, ${\cal U}$ and ${\cal O}$ represent mask token, IoU token, and semantic token, respectively.
-In the column of the IoU head, $F$ means to freeze the original IoU head and train a parallel one and $T$ means tuning the original IoU head.
 Ablation on PWD-Net.
-We compare PWD-Net with variants that replace some tokens with a full-zero placeholder.
-We also compare with two designs, tuning the IoU head directly or learning a parallel IoU head.
+We compare PWD-Net to the variants that replace some tokens with a full-zero placeholder.
+We also consider two designs, directly tuning the IoU head or learning a parallel IoU head.
 As shown in  Tab.
-8, all three tokens, i.e.
+8, all the three tokens, i.e.
 Mask Token ${\cal M}$, IoU Token ${\cal U}$, and Semantic Token ${\cal O}$, contribute to the final result.
 Particularly, once ${\cal M}$ is removed, the AP drops by 40.0%, which is a catastrophic decline.
-This degradation suggests that the Mask Token contains the essential shape-aware feature that is essential for the part-whole discrimination task.
-Notably, the AP drops by 2.8% when we tune the pre-trained IoU Head, suggesting that it might overfit the few labeled images.
-We believe by freezing the SAM’s IoU head, the PWD-Net can benefit more from the shape-aware knowledge learned on massive segmentation data.
-Figure 4: Qualitative comparisons between our proposed Crowd-SAM (a) and De-FRCN (b).
-We also plot the generated masks and GT boxes (c) and our prompts filtering results (d), where red points are kept and gray points are filtered.
+This degradation suggests that the mask token contains the shape-aware feature that is essential for the part-whole discrimination task.
+Notably, the AP drops by 2.8% when we tune the pre-trained IoU Head, suggesting that it is prone to overfit the few labeled images.
+By freezing the IoU head of SAM, PWD-Net can benefit more from the shape-aware knowledge learned from massive segmentation data.
+Figure 4: Qualitative comparison between Crowd-SAM (a) and De-FRCN (b).
+Crowd-SAM predictions are more accurate especially in the boundaries of persons.
+We also plot the GT boxes (blue rectangles) and the generated masks (yellow regions), which are of high quality (c).
+In (d), we plot our prompt filtering results, where preserved prompts (red points) are much fewer than the removed ones (gray points).
+Zoom in for a better view.
 5 Conclusion
-This paper introduces Crowd-SAM, a SAM-based framework for object detection and segmentation in crowded scenes, designed to streamline the annotation process.
+This paper proposes Crowd-SAM, a SAM-based framework, for object detection and segmentation in crowded scenes, designed to streamline the annotation process.
 For each image, Crowd-SAM generates dense prompts for high recall and uses EPS to prune redundant prompts.
-To achieve accurate detection in occlusion cases, Crowd-SAM employs PWD-Net that leverages several informative tokens to select the best-fitting mask.
-Combined with the proposed methods, Crowd-SAM achieves 78.4 AP on Crowd Human, comparable to full-supervised detectors, validating that object detection in crowded scenes can benefit from foundation models like SAM in data efficiency.
+To achieve accurate detection in occlusion cases, Crowd-SAM employs PWD-Net which leverages several informative tokens to select the masks that best fit.
+Combined with the proposed modules, Crowd-SAM achieves 78.4% AP on CrowdHuman, comparable to full-supervised detectors, validating that object detection in crowded scenes can benefit from foundation models like SAM with data efficiency.
 Acknowledgements